### v1

@@ 3.1 Color-Magnitude Diagrams
The left panels of Figure 1 (a) and (b) show the color-magnitude diagrams (CMDs) for Cen I and Eri IV which include stars within one half-light radius (as determined in Section 3.5, see Table 4). Plotted in red are 13.0 Gyr, metal-poor PARSEC isochrones (Bressan et al., 2012). This corresponds to the best-fit isochrone as determined in Section 3.2. The magenta error bars show the mean photometric errors in color and magnitude as determined by the artificial star tests (see Section 2). These errors are plotted at an arbitrary color for convenience. Open blue diamonds are potential horizontal branch (HB) stars within the half-light radius and green triangles are potential HB stars within the FoV: a total of eleven in Cen I and seven in Eri IV. Newly discovered HB candidates are presented in Tables 5 and 6. In the CMD for Cen I we show two RR Lyrae stars from Martínez-Vázquez et al. (2021) as gray triangles; the third star from that study lies outside our FoV. The filled triangle lies within our half-light radius, while the other RR Lyrae star is within the FoV.
@@ 3.2 Distance
We also derive a distance modulus for both dwarfs using their potential HB stars within the FoV. This is 11 stars for Cen I and 7 stars for Eri IV (see Figure 1). We fit the M92 globular cluster HB fiducial (Bernard et al., 2014) to our HB star candidates by minimizing the sum of squares of the difference between the data and the fiducial. We measure the magnitude offset a given HB star lies from the fiducial, shift the distance modulus of the fiducial by 0.025 mag, and remeasure the offset. This process is repeated through two magnitudes. The shift that minimizes the sum of squares is adopted as the distance modulus. We opt to use empirical HB fiducial tracks rather than HB isochrones as HB isochrones are historically more difficult to model and have a certain level of uncertainty associated with them (e.g. Pietrinferni et al., 2004). We use M92 ( mag, Bernard et al. 2014) specifically because it is one of the most ancient, metal-poor, and well-studied globular clusters.
@@ 3.3 Structural Properties
To determine the structural properties of Cen I and Eri IV, we use a maximum likelihood analysis described in Sand et al. (2009) (see also Martin et al., 2008). In short, the routine fits an exponential profile to the 2D distribution of stars associated with each UFD. We select stars consistent within our 90% completeness measurement and a PARSEC isochrone in color-magnitude space after accounting for photometric uncertainties. The structural parameters from the discovery papers (Mau et al., 2020; Cerny et al., 2021) are adopted as the input parameters for the initial analysis. The data are bootstrap resampled 1000 times; the structural parameters are recalculated for each such resampling which determines the uncertainty on the measurement. The resulting structural parameters are shown in Table 4.
@@ 3.5 Extended Structure Search
To apply the matched-filter technique to our data, we use the same simulated stars as used in Section 3.4 as the signal CMD (i.e., stars simulated to match an old, metal-poor PARSEC isochrone, accounting for photometric errors and magnitudes brighter than the 90% completeness threshold). Simulated data is preferable to the observed data which can be sparsely populated and contaminated with background/foreground objects (Sand et al., 2012). For the background CMD, we use the real stars well outside the half-light radius of both UFDs. The matched-filter stellar density maps are shown in Figure 2; both maps have been spatially binned to a pixel size of 25″ and smoothed with a Gaussian width 1.0 times the pixel size. The background and variance of the smoothed maps are determined with IDL’s MMM routine. The main body of each satellite is clearly visible in each map. The white arrows in Figure 2 represent the direction to the Galactic center, and the magenta arrows denote the solar-reflex-corrected proper motions from Heiger et al. (2024). The spectroscopic member stars observed by Heiger et al. (2024) are shown in red (RGB) and blue (HB), and the potential member stars discussed in Section 3.6 are shown in green.
@@ 3.6 Potential Member Stars in Gaia
We also explore the entire FoV by cross-matching stars that agree with our CMD selection used in Section 3.3 with Gaia EDR3. We keep stars that satisfy the following criteria: ruwe , astrometric_excess_noise , and stars which are consistent with zero parallax (). Furthermore, we exclude stars which are inconsistent with the proper motions of the stars presented in Heiger et al. (2024). This flexible selection method yields four additional stars in Cen I and eight additional stars in Eri IV. The potential new member stars in Cen I and Eri IV are listed in Tables 5 and 6, respectively. They are ideal targets for a future spectroscopic study. Their spatial positions are shown in Figure 2 in green. Two of Eri IV’s newly identified potential member stars are situated near the extended feature.
@@ 4.1 Cen I
The deep imaging of Cen I shows a well-populated MS, an easily identifiable RGB, eleven potential HB candidates within the FoV (four within one ), and a handful of blue stragglers (Figure 1, left). The stellar population is consistent with an age of Gyr and a metallicity of . We find the distance to Cen I is kpc ( mag); this is in good agreement with the results reported in Mau et al. (2020) ( kpc, mag). Our measurement also agrees well with Martínez-Vázquez et al. (2021) who derived the distance using three RR Lyrae stars. They found kpc, mag with systematic errors of 4 kpc and 0.07 mag. We show two of the RR Lyrae stars in Figure 1, the third star was outside our FoV.
@@ 4.2 Eri IV
The CMD of Eri IV shows a MS with a fair amount of color scatter (due to photometric uncertainties), but there exists a clear MS-turnoff. The RGB is less populated compared to Cen I, which aligns with our findings that Cen I is brighter. We find seven potential HB stars within the FoV (two within one ) and a number of potential blue stragglers (Figure 1). The stellar population of Eri IV is consistent with an age of Gyr and a metallicity of . We find Eri IV to be marginally closer than initially reported, but consistent within the joint uncertainty: kpc (this work), compared to kpc from Cerny et al. (2021).
@@ 5 Summary and Conclusion
Our work helps to clear up some of the open questions surrounding the discovery photometry and the spectroscopic data: both Cen I and Eri IV exhibited hints of extended features in the discovery data, but spectroscopic analysis suggested these overdensities were not the result of tidal disruption. Our photometry shows no compelling evidence that Cen I is disrupting. Eri IV, on the other hand, hosts an extended structure to the northeast of its central body (Figure 2). This structure may be the result of tidal disruption in the recent past, although we can’t rule out the possibility that this region arose from a merger or it is a satellite itself. It seems as though Eri IV will remain an interesting UFD moving forward. This work highlights the importance of deep imaging in conjunction with spectroscopy to specifically explore the extended structures, ultimately aiding in the understanding of the true nature of ultra-faint satellites.
@@ 6 Acknowledgments
QOC acknowledges support from the Dartmouth Fellowship. DJS and the Arizona team acknowledge support from NSF grant AST-2205863. QOC thanks Ryan Hickox and Elisabeth Newton for useful discussions.

### v2

@@ 3.1 Color-Magnitude Diagrams
The left panels of Figure 1 (a) and (b) show the color-magnitude diagrams (CMDs) for Cen I and Eri IV which include stars within one half-light radius (as determined in Section 3.3, see Table 4). Plotted in red are 13.0 Gyr, PARSEC isochrones (Bressan et al., 2012). This corresponds to the best-fit isochrone as determined in Section 3.2. The magenta error bars show the mean photometric errors in color and magnitude as determined by the artificial star tests (see Section 2). These errors are plotted at an arbitrary color for convenience. Open blue diamonds are potential horizontal branch (HB) stars within the half-light radius and green triangles are potential HB stars within the FoV: a total of eleven in Cen I and seven in Eri IV. Newly discovered HB candidates are presented in Tables 5 and 6. In the CMD for Cen I we show two RR Lyrae stars from Martínez-Vázquez et al. (2021) as purple triangles; the third star from that study lies outside our FoV. The filled triangle lies within our half-light radius, while the other RR Lyrae star is within the FoV.
@@ 3.2 Distance
We also derive a distance modulus for both dwarfs using their potential HB stars within the FoV. This is 11 stars for Cen I and 7 stars for Eri IV (see Figure 1). The number of potential HB stars in Cen I and Eri IV is comparable to other well-studied systems (e.g. Sand et al., 2010; Koposov et al., 2015; Muñoz et al., 2018; Homma et al., 2019). We fit the M92 globular cluster HB fiducial (Bernard et al., 2014) to our HB star candidates by minimizing the sum of squares of the difference between the data and the fiducial. We measure the magnitude offset a given HB star lies from the fiducial, shift the distance modulus of the fiducial by 0.025 mag, and remeasure the offset. This process is repeated through two magnitudes. The shift that minimizes the sum of squares is adopted as the distance modulus. We opt to use empirical HB fiducial tracks rather than HB isochrones as HB isochrones are historically more difficult to model and have a certain level of uncertainty associated with them (e.g. Pietrinferni et al., 2004). We use M92 ( mag, Bernard et al. 2014) specifically because it is one of the most ancient, metal-poor, and well-studied globular clusters.
@@
Old, low-metallicity isochrones exhibit minimal variation based on the chosen age and metallicity. Separate libraries, such as the Dartmouth Stellar Evolution Database (DSEP; Dotter et al., 2008), closely resemble PARSEC isochrones in this regime. Using DSEP isochrones ( and ) we derive distances consistent with the PARSEC models within the uncertainties quoted in Table 4; we show one DSEP isochrone in Figure 1 for reference. Ultimately, we choose the PARSEC models to remain consistent with the discovery papers.
@@ 3.3 Structural Properties
To determine the structural properties of Cen I and Eri IV, we use a maximum likelihood analysis described in Sand et al. (2009) (see also Martin et al., 2008). In short, the routine fits an exponential profile to the 2D distribution of stars associated with each UFD. We select stars consistent within our 90% completeness measurement and a PARSEC isochrone in color-magnitude space after accounting for photometric uncertainties. This is 812 stars for Cen I and 243 stars for Eri IV. The structural parameters from the discovery papers (Mau et al., 2020; Cerny et al., 2021) are adopted as the input parameters for the initial analysis. The data are bootstrap resampled 1000 times; the structural parameters are recalculated for each such resampling which determines the uncertainty on the measurement. The resulting structural parameters are shown in Table 4.
@@ 3.5 Extended Structure Search
To apply the matched-filter technique to our data, we use the same simulated stars as used in Section 3.4 as the signal CMD (i.e., stars simulated to match an old, metal-poor PARSEC isochrone, accounting for photometric errors and magnitudes brighter than the 90% completeness threshold). Simulated data is preferable to the observed data which can be sparsely populated and contaminated with background/foreground objects (Sand et al., 2012). For the background CMD, we use the real stars well outside the half-light radius of both UFDs. The matched-filter stellar density maps are shown in Figure 2; both maps have been spatially binned to a pixel size of 25″ and smoothed with a Gaussian width 1.0 times the pixel size. The background and variance of the smoothed maps are determined with IDL’s MMM routine. The main body of each satellite is clearly visible in each map. The white arrows in Figure 2 represent the direction to the Galactic center, and the magenta arrows denote the solar-reflex-corrected proper motions from Heiger et al. (2024). The spectroscopic member stars observed by Heiger et al. (2024) are shown in red (RGB) and blue (HB), and the potential member stars discussed in Section 3.6 are shown in orange.
@@ 3.6 Potential Member Stars in Gaia
We also explore the entire FoV by cross-matching stars that agree with our CMD selection used in Section 3.3 with Gaia EDR3. We keep stars that satisfy the following criteria: ruwe , astrometric_excess_noise , and stars which are consistent with zero parallax (). Furthermore, we exclude stars which are inconsistent with the proper motions of the stars presented in Heiger et al. (2024). This flexible selection method yields four additional stars in Cen I and eight additional stars in Eri IV. The potential new member stars in Cen I and Eri IV are listed in Tables 5 and 6, respectively. They are ideal targets for a future spectroscopic study. Their spatial positions are shown in Figure 2 in orange. Two of Eri IV’s newly identified potential member stars are situated near the extended feature.
@@ 4.1 Cen I
The deep imaging of Cen I shows a well-populated MS, an easily identifiable RGB, eleven potential HB candidates within the FoV (four within one ), and a handful of blue stragglers (Figure 1, left). The stellar population is consistent with an age of Gyr and a metallicity of . We find the distance to Cen I is kpc ( mag); this is in good agreement with the results reported in Mau et al. (2020) ( kpc, mag). Our measurement also agrees well with Martínez-Vázquez et al. (2021) who derived the distance using three RR Lyrae stars. They found kpc, mag with systematic errors of 4 kpc and 0.07 mag. We show two of the RR Lyrae stars in Figure 1, the third star was outside our FoV. Three RR Lyrae stars is consistent with the expected number ( stars) for UFDs of similar magnitudes (Martínez-Vázquez et al., 2019, 2021).
@@ 4.2 Eri IV
The CMD of Eri IV shows a MS with a fair amount of color scatter (due to photometric uncertainties), but there exists a clear MS-turnoff. The RGB is less populated compared to Cen I, which aligns with our findings that Cen I is brighter. We find seven potential HB stars within the FoV (two within one ) and a number of potential blue stragglers (Figure 1). We find one potential RR Lyrae star (Table 6) which appears in the Gaia DR3 variability catalog (Eyer et al., 2023), but is classified as an eclipsing binary. A system with the luminosity of Eri IV is expected to have few, if any, RR Lyrae stars (Martínez-Vázquez et al., 2019). The stellar population of Eri IV is consistent with an age of Gyr and a metallicity of . We find Eri IV to be marginally closer than initially reported, but consistent within the joint uncertainty: kpc (this work), compared to kpc from Cerny et al. (2021).
@@ 5 Summary and Conclusion
Our work helps to clear up some of the open questions surrounding the discovery photometry and the spectroscopic data: both Cen I and Eri IV exhibited hints of extended features in the discovery data, but spectroscopic analysis suggested these overdensities were not the result of tidal disruption. Our photometry shows no compelling evidence that Cen I is disrupting. Eri IV, on the other hand, hosts an extended structure to the northeast of its central body (Figure 2). This structure may be the result of tidal disruption in the recent past, although we can’t rule out the possibility that this region arose from a merger or it is a satellite itself. It seems as though Eri IV will remain an interesting UFD moving forward; the true nature of the extended structure may be understood through a combination of deeper and wider-field imaging data and detailed dynamical modeling. This work highlights the importance of deep imaging in conjunction with spectroscopy to specifically explore the extended structures, ultimately aiding in the understanding of the true nature of ultra-faint satellites.
@@ 6 Acknowledgments
We thank the anonymous reviewer whose comments and suggestions improved the content of this paper. QOC acknowledges support from the Dartmouth Fellowship. DJS and the Arizona team acknowledge support from NSF grant AST-2205863. QOC thanks Ryan Hickox and Elisabeth Newton for useful discussions.